OpenStudy (anonymous):

MEDAL GIVEN - Mr. and Mrs. Dawson's son Josh weighed 28 pounds on his second birthday. On his eighth birthday, Josh weighed 56 pounds. What is the percentage increase in Josh's weight from his second birthday until his eighth birthday? A. 28% B. 50% C. 84% D. 100%

OpenStudy (vanessagilbert1023):

I would say D ..

OpenStudy (vanessagilbert1023):

56-28 divided by 28 is equal to 28/28 X 100

jimthompson5910 (jim_thompson5910):

The change in weight is 56 - 28 = 28 lbs Divide this change by the original weight: 28/28 = 1 Then multiply by 100 to get the percentage: 1*100 = 100% So this is a 100% increase (basically the weight doubled)

OpenStudy (anonymous):

@vanessagilbert1023 @jim_thompson5910 thanks guys (: